Minutes PSB Upgrade WG Meeting 214, 8th of May 2018
Participants: Wolfgang Bartmann, Aurelio Berjillos Barranco, Thomas William Birtwistle, Jean-Baptiste Bonammy, Louis De Mallac, Gian Piero Di Giovanni, Alessandro Floriduz, Vincenzo Forte, Robert Froeschl, Matthias Haase, B. Mikulec, Gabor Petrika, Serge Pittet, Jocelyn Tan, Wim Weterings.
Indico event: https://indico.cern.ch/event/681817/
Agenda:
1. Approval of minutes
The minutes of the LIU-PSB WG #213
meeting were approved.
2. Communications
- LIU Project Team Meetings:
- Last meeting we reviewed the analysis done for each WP to verify which money could be saved and which would be the impact on performance.
- Thanks a lot to all WP-holders who helped preparing the summary
- Saving have been identified from TE-ABT (extraction kicker) and EN-HE (transport). The saved money will be allocated in a special budget in the management workpackage to avoid changing the cost-to-completion of the LIU-PSB project presented at the last cost&schedule review. In future, if any of the wokpackage identifies any saving, the money we will be stored in the same work-package.
- Other news is that Linac4 team report at the LIU Beam Performance Meeting about the beam quality run.
- Linac4 will stop on Friday 11th May and enter a long technical stop until 17th August. The length of the stop could not be reduced and it is mostly driven by the work to be done on the RF systems.
3. EVM PSB Update (S. Prodon)
- The cost variance (CV = EV-AC) is positive 1,794 kCHF and the schedule variance (SV = PV-EV) slightly negative -0.39 months:
- As during the Cost&Schedule Review in 2018 there were couple of large invoices (RF and cabling) not yet paid and which account for the large positive variance.
- The schedule variance is indeed small but the reason is that TE-EPC has a large positive schedule variance (1,053 kCHF) covering almost all the negative variances.
- The highlighted workpackages which are currently late are BE-BI (SV = -96 kCHF), BE-RF (SV = -127 kCHF), EN-CV (SV = -220 kCHF), TE-ABT (SV = -181 kCHF), TE-MSC (SV = -297 kCHF), TE-VSC (SV = -308 kCHF).
- S. Prodon mentioned that the commitment curve shows a only as slow increase, so it would be important to keep an eye on the commitments in the project in the next weeks.
- As a general remark, it is very important that each WU-holder updates their progress regularly to provide a clear and steady picture of the advancement of the project.
4. Report on the ECRs for LS2 (T. Birtwistle)
Multiple groups:
- PSB transfer lines (BT, BTM, BTP): magnet modifications, beam instrumentations, beam stopper, vacuum changes. Work is advancing well and W. Bartmann is in contact with T. Birtwistle and integration to sort out the details.
- New BPM on the BI line: no news → W. Bartmann mentioned that C. Hessler was only waiting for some integration drawing. He will double check with him.
- Modifications to the PSB extraction main unit: work ongoing. A first draft expected for mid-May.
TE-ABT:
- Very good progress.
- ECR about 'replacement of the extraction kicker KFA14L1' is under approval. Waiting for impedance assessment.
- G.P. Di Giovann*i asked if the aperture is going to change. *W. Weterings mentioned that it should not be the case, but it will be explicitly mentioned in the released version.
- Last remaining ECR about 'replacing the kicker BT.KFA10' is in work. Draft is ready but some clarification about the attached BTV are needed.
TE-MSC:
- Last remaining ECR is about the 'upgrade of the BHZ main unit bending magnets'. Will take several weeks, integration work is ongoing.
TE-EPC:
- Most of them under approval.
- Potential impact on the septum transformers. Experts working on it. T. Birtwistle will contact J.M. Cravero.
BE-BI:
- The last ECR is about the 'new BLMs installation in the LT, LTB and BI Lines for H- exchange monitoring'. A. Lombardi agreed to discuss with J.P. Corso to define the exact location.
BE-RF:
EN-STI:
- The ECR will be sent very soon. Integration work has to be completed.
EN-CV:
- The ECR of the consolidation of the PSB cooling systems are in draft and the integration is missing.
- G. Petrika mentioned that the ECR about the new ventilation system will only come late in time as EN-CV wants to focus on completing the work for the cooling plants first.
EN-ACE:
- The ECR are under approval.
- Concerning the latest circulation of the ECR on the partial dismantling of the Linac2, EN-ACE will look after the orphan equipment trying to identify who could remove them and the associated budget.
5. AOB
- R. Froeschl will officially step down as RP contact person for LIU-PSB. F. Pozzi will be the new contact person. R. Froeschl will remain deputy for the next month or so to smooth the transition.
- G.P. Di Giovanni thanked R. Froeschl for this dedication to the project since the beginning and wished success for his future endeavors.
Next LIU-PSB meeting: 22nd of May 2018 on the status of the H0/H- current monitor and SEM-grid for LBE line (F. Roncarolo) and planning for LS2 (D. Hay).
There are minutes attached to this event.
Show them.